# Relativistic stars in $f(Q)$-gravity: Exact analytic solution for the power-law case $f(Q)=Q+b\>Q^{\nu}$

### v2

@@ Abstract
We investigate static spherically symmetric spacetimes within the framework of symmetric teleparallel $f(Q)$ gravity in order to describe relativistic stars. We adopt a specific ansatz for the background geometry corresponding to a singularity-free space-time. We obtain an expression for the connection, which allows the derivation of solutions for any $f(Q)$ theory in this context. Our approach aims to address a recurring error appearing in the literature, where even when a connection compatible with spherical symmetry is adopted, the field equation for the connection is systematically omitted and not checked if it is satisfied. For the stellar configuration, we concentrate on the power-law model $f(Q)=Q+\alpha Q_{0}\left(\frac{Q}{Q_{0}}\right)^{\nu}$. The de Sitter-Schwarzschild geometry naturally emerges as an attractor beyond a certain radius, we thus utilize it as the external solution beyond the boundary of the star. We perform a detailed investigation of the physical characteristics of the interior solution, explicitly determining the mass function, analyzing the resulting gravitational fluid properties and deriving the angular and radial speed of sound.
Stars; Static Spherically Symmetric; Symmetric teleparallel $f(Q)$-gravity
@@ I Introduction
Possible quantum corrections to GR are sometiems expressed through modifications of the Einstein-Hilbert Action by the addition of higher-order curvature invariants [12, 13, 14, 15]. The Starobinsky model of inflation [16] belongs to this family of theories; an $R^{2}$ term is introduced in order to drive the dynamics that explain the early-time acceleration of the universe. The general family of modified theories in which the Starobinsky model belongs to is the $f\left(R\right)$-gravity [17]. These theories are also related, through conformal mappings, to scalar tensor theories like e.g. the Brans-Dicke gravity [18].
@@
There are several studies in the literature that investigate static, spherically symmetric spacetimes in $f\left(Q\right)$-gravity. Exact solutions describing black holes have been presented in [47, 48, 49, 50, 51, 52], while black holes in nonmetricity scalar-tensor theories were studied in [50, 53]. The latter, is a more general theory which includes $f\left(Q\right)$-gravity as a limiting case. The setting is similar to that which relates Brans-Dicke to $f\left(R\right)$-gravity. In [53], an interesting result was found: namely, there are no static spherically symmetric black hole solutions in $f\left(Q\right)$-gravity when the two metric coefficients, $g_{tt}$ and $g_{rr}$, are constrained by $g_{rr}g_{tt}=1$. Moreover, the use of the coincidence gauge for static, spherically symmetric spacetimes was discussed in [54], and possible constraints from solar system tests in the theory are discussed in [55].
As has been mentioned in [56], the timeline of the investigation of the properties of relativistic and compact stars, initiated in 1916 with Schwarzschild’s solution to Einstein’s field equations; an analysis that has been fundamental to astrophysics. The solution may be interpreted as the gravitational field generated by a body characterized by spherical symmetry [57]. That work boosted subsequent studies on stellar modelling and internal composition of astronomical objects characterized by ultra dense matter content. Under those extreme conditions that cannot reproduced on terrestrial experiments, the Newtonian description becomes inadequate, and thus a full general relativistic treatment is required. Thanks to their unique properties, relativistic stars comprise ideal cosmic laboratories capable of constraining new physics and alternative theories of gravity.
When studying stars the authors usually focus their attention on astrophysical objects made of matter viewed as a perfect fluid, where there is a unique pressure along all spatial dimensions, $p_{r}=p_{\theta}=p_{\phi}$. However, celestial bodies are not necessarily made of isotropic fluids. In fact, under certain conditions matter content of compact objects may become anisotropic. The review article of Ruderman [58] long time ago mentioned such a possibility: there the author makes for the first time the observation that in very dense media anisotropies can be generated by interactions between relativistic particles. A comprehensive study of fluid spheres made of anisotropic matter was presented in [59] in Newtonian as well as relativistic gravity. In the context of Einstein’s General Relativity, anisotropic fluids have been extensively studied, with particular emphasis on stellar modelling and their impact on structural properties of compact objects [60, 61].
@@ III.1 Relativistic stars within General Relativity
When seeking interior solutions ($0\leq r\leq R$, $R$ being the radius of the star), the most general form of a static, spherically symmetric geometry in Schwarzschild-like coordinates, $\{t,r,\theta,\phi\}$, is given by
@@
$p_{r}(R)=0,\;\;\;\;m(R)=M,\;\;\;\;e^{A(R)}=1-2\frac{M}{R},$ (25)
with $M$ being the stellar mass, taking into account that the exterior vacuum solution ($r>R$) is given by the Schwarzschild geometry [57]
@@
$A(r)=\ln\left(1-\frac{2M}{R}\right)+2\int^{r}_{R}dx\>\frac{m(x)+4\pi x^{3}p_{r}(x)}{x^{2}\left(1-2m(x)/x\right)}.$ (27)
@@ III.2 Exact analytic solution within $f(Q)$ gravity
The theory is characterized by two basic fields, the metric and the connection. For the former we adopt the static and spherically symmetric line element as seen in Eq. (17). In what regards the connection, it is well known [74] that for a flat and symmetric connection there can always be found a coordinate system where its components are zero, i.e. $\Gamma_{~\mu\nu}^{\alpha}=0$. This is referred to the literature as the coincident gauge. However, we need to be careful when we impose such a condition. When assuming a particular class of metrics, like the one we read from (17), we need to consider that we have already fixed the gauge by choosing a particular coordinate system. The latter may happen to be incompatible with having $\Gamma_{~\mu\nu}^{\alpha}=0$.
In these cases, a guide in order to write connections which are compatible with the field equations, is given by requiring that the connection shares the symmetries of the metric, i.e. the isometries. In other words we demand that the Killing fields of the metric, satisfying $\mathcal{L}_{\xi}g_{\mu\nu}=0$, where $\mathcal{L}$ is the Lie-derivative, leave invariant the connection $\mathcal{L}_{\xi}\Gamma_{~\mu\nu}^{\alpha}=0$ as well. The resulting connections can play an important role in the dynamics introducing additional degrees of freedom [37]. For static, spherically symmetric spacetimes, there are two general families of connections with this property [75, 47]. The first leads to off-diagonal components in the field equations (10) for the gravitational part. These off-diagonal terms can only be eliminated by setting Q=Q=const. or $f(Q)\sim Q$. Alternatively, one can keep the off-diagonal terms and consider a non-diagonal energy momentum tensor for the matter. For the second family of connections the relative non-diagonal terms can be removed without requiring a constraint on the theory $f(Q)$ or on $Q$ itself. It is this second connection which we choose in this work. Its non-zero components are [75, 47, 50]:
@@
Non-metricity theories are based on two basic fields, the metric $g_{\mu\nu}$ and the connection $\Gamma_{\;\mu\nu}^{\kappa}$. The set of components (28) forms the initial ansatz for the connection $\Gamma_{\;\mu\nu}^{\kappa}$ in the same manner line element (17) constitutes the ansatz for the metric. For this choice of a connection, and for metric (17), the off-diagonal terms of the field equations (10) are eliminated if we further choose $c_{1}=c_{2}=0$.
@@
satisfies equation (30) irrespectively of the $f(Q)$ function and of the $A(r)$ and $B(r)$ appearing in the metric. This connection has been also utilized in the context of scalar non-metricity to derive black hole solutions [22].
A brief comment is in order at this point. Many works in the literature, attempting to study stellar configurations in $f(Q)$ gravity, either incorporate an overrestricting procedure or include the serious error of disregarding the equation of the connection. In some cases, the coincident gauge is adopted, which however is incompatible with line element (17) in spherical coordinates and thus overrestricts the equations; leading only to solutions of the linear $f(Q)\sim Q$ theory, which is however equivalent to GR [76, 77, 78, 79, 80]. In other cases, even though a connection compatible with spherical symmetry is considered, it is usually not checked whether the equation of motion for the connection (12) is satisfied. To make this problem more explicit; a common choice of a connection used in the literature [81, 82, 83, 84, 85] is the following
@@
Remember that $Q$ here is given by (34), and for a given $f(Q)$ theory this is going to be purely a relation between the functions of the metric, $A$ and $B$. The above equation is trivialized if Q=Q=const. or $f(Q)\sim Q$, which, as is known, describe the same solutions as in the case of GR with or without cosmological constant. Thus, in the generic setting of a non-trivial $f(Q)$ theory, it is not sufficient for the description of the star solution to just solve algebraically (29) with respect to the energy density and the pressures. The constraint set by the field equation of the connection needs to be satisfied. Note that it is not enough to satisfy the analogue of the Tolman-Oppenheimer-Volkoff (TOV) equation. The $\nabla_{\mu}T^{\mu\nu}=0$ condition on the matter content together with the field equations for the metric (29) are not sufficient conditions. This is because, as we previously mentioned, the covariant derivative of the gravitational tensor $E_{\mu\nu}$ of Eq. (13), unlike what happens in GR, is not identically zero. The $\nabla_{\mu}E^{\mu\nu}=0$ equation holds only if the equation of the connection (30) is satisfied [71]. In order for the latter to happen for a theory dynamically different from GR, we need to find an appropriate $\gamma_{2}(r)$ function that solves the connection equation.
In our case, the $\gamma_{2}(r)$ we derived in (32) serves to satisfy the field equation for the connection (30), and thus the constraint in the metric functions is overall avoided. What is more, (32) satisfies the relevant equation for any $f(Q)$ theory. In this respect, the $\gamma_{2,\text{int}}=re^{-\frac{B}{2}}$ forms an excellent choice for the connection in what regards the interior solution for the star. For the exterior vacuum solution we adopt a different path, which we describe in detail in the next section. It consists of adopting an exterior solution, $\gamma_{2,\text{ext}}$, for which the non-metricity scalar becomes constant. Thus, we achieve for the inside, a consistent description of modified gravitational dynamics due to a dynamical non-metricity, while in the exterior a known vacuum solution compatible with GR plus a cosmological constant. In the same manner that one connects two distinct solutions for the interior (with matter) and for the exterior (vacuum) for the metric, we additionally connect here the two distinct solutions for the other basic field of the theory, the connection.
@@ III.3 Exterior solution
where the effective cosmological constant $\Lambda_{\text{eff}}$ is given by (15).
@@ IV The model under consideration
where $\alpha$, $Q_{0}$ and $\nu$ are constants. The ansatz we follow for the metric functions $A$ and $B$ is the following [86]
@@
with $C\in\mathbb{R}$ and $r_{1}$, $r_{2}$ positive constants. The corresponding geometry describes a singular-free physical solution.
@@
Using Eq. (44) in the expressions (41)-(43), and considering theories $f(Q)$ for which $\nu>1$ in (38), it is straightforward to demonstrate that the relevant physical quantities assume a finite value at $r=0$. The nonmetricity scalar itself becomes zero at $r\rightarrow 0$. In the Taylor expansion, there always exists a zero-th order term, while the rest that follow are in positive powers of $r$. Specifically, for $\nu\geq 2$ the leading $r$-dependent terms are quadratic:
@@
$c_{s,\text{radial}}^{2}=\frac{p_{r}^{\prime}}{\rho^{\prime}},\quad c_{s,\text{tangential}}^{2}=\frac{p_{t}^{\prime}}{\rho^{\prime}}$ (48)
we obtain at the limits $r\rightarrow 0$
$\lim\limits_{r\to 0}c_{s,\text{radial}}^{2}=\frac{\left(4r_{2}^{2}-r_{1}^{2}\right)\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+4\alpha r_{2}^{2}\right)}{5r_{1}^{2}\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+8\alpha r_{2}^{2}\right)},$ (49)
$\lim\limits_{r\to 0}c_{s,\text{tangential}}^{2}=\frac{4\alpha\left(r_{1}^{4}-6r_{1}^{2}r_{2}^{2}+8r_{2}^{4}\right)-2Q_{0}r_{2}^{2}\left(r_{1}^{4}-3r_{1}^{2}r_{2}^{2}+r_{2}^{4}\right)}{5r_{1}^{2}\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+8\alpha r_{2}^{2}\right)}$ (50)
@@
$\lim\limits_{r\to 0}c_{s,\text{radial}}^{2}=\frac{4r_{2}^{2}}{5r_{1}^{2}}-\frac{1}{5},$ (51)
$\lim\limits_{r\to 0}c_{s,\text{tangential}}^{2}=\frac{2\left(3r_{1}^{2}r_{2}^{2}-r_{1}^{4}-r_{2}^{4}\right)}{5r_{1}^{4}}.$ (52)

### v3

@@ Abstract
We investigate static spherically symmetric spacetimes within the framework of symmetric teleparallel $f(Q)$ gravity in order to describe relativistic stars. We adopt a specific ansatz for the background geometry corresponding to a singularity-free spacetime. We obtain an expression for the connection, which allows the derivation of solutions for any $f(Q)$ theory in this context. Our approach aims to address a recurring error appearing in the literature, where even when a connection compatible with spherical symmetry is adopted, the field equation for the connection is systematically omitted and not checked if it is satisfied. For the stellar configuration, we concentrate on the power-law model $f(Q)=Q+\alpha Q_{0}\left(\frac{Q}{Q_{0}}\right)^{\nu}$. The de Sitter-Schwarzschild geometry naturally emerges as an attractor beyond a certain radius, we thus utilize it as the external solution beyond the boundary of the star. We perform a detailed investigation of the physical characteristics of the interior solution, explicitly determining the mass function, analyzing the resulting gravitational fluid properties and deriving the angular and radial speed of sound.
Relativistic Stars; Static Spherically Symmetric; Exact Solutions; Symmetric teleparallel $f(Q)$-gravity
@@ I Introduction
Possible quantum corrections to GR are sometimes expressed through modifications of the Einstein-Hilbert Action by the addition of higher-order curvature invariants [12, 13, 14, 15]. The Starobinsky model of inflation [16] belongs to this family of theories; an $R^{2}$ term is introduced in order to drive the dynamics that explain the early-time acceleration of the universe. The general family of modified theories in which the Starobinsky model belongs to is the $f\left(R\right)$-gravity [17]. These theories are also related, through conformal mappings, to scalar tensor theories like e.g. the Brans-Dicke gravity [18].
@@
There are several studies in the literature that investigate static, spherically symmetric spacetimes in $f\left(Q\right)$-gravity. Exact solutions describing black holes have been presented in [47, 48, 49, 50, 51, 52, 53], while black holes in nonmetricity scalar-tensor theories were studied in [50, 54]. The latter, is a more general theory which includes $f\left(Q\right)$-gravity as a limiting case. The setting is similar to that which relates Brans-Dicke to $f\left(R\right)$-gravity. In [54], an interesting result was found: namely, there are no static spherically symmetric black hole solutions in $f\left(Q\right)$-gravity when the two metric coefficients, $g_{tt}$ and $g_{rr}$, are constrained by $g_{rr}g_{tt}=-1$. Moreover, the use of the coincidence gauge for static, spherically symmetric spacetimes was discussed in [55], and possible constraints from solar system tests in the theory are discussed in [56]. Recently neutron stars solutions in $f(Q)$ theory have been presented in [57], and a stellar solution in the context of scalar nonmetricity in [58].
As has been mentioned in [59], the timeline of the investigation of the properties of relativistic and compact stars, initiated in 1916 with Schwarzschild’s solution to Einstein’s field equations; an analysis that has been fundamental to astrophysics. The solution may be interpreted as the gravitational field generated by a body characterized by spherical symmetry [60]. That work boosted subsequent studies on stellar modelling and internal composition of astronomical objects characterized by ultra-dense matter content. Under those extreme conditions that cannot reproduced on terrestrial experiments, the Newtonian description becomes inadequate, and thus a full general relativistic treatment is required. Thanks to their unique properties, relativistic stars comprise ideal cosmic laboratories capable of constraining new physics and alternative theories of gravity.
When studying stars the authors usually focus their attention on astrophysical objects made of matter viewed as a perfect fluid, where there is a unique pressure along all spatial dimensions, $p_{r}=p_{\theta}=p_{\phi}$. However, celestial bodies are not necessarily made of isotropic fluids. In fact, under certain conditions matter content of compact objects may become anisotropic. The review article of Ruderman [61] long time ago mentioned such a possibility: there the author makes for the first time the observation that in very dense media anisotropies can be generated by interactions between relativistic particles. A comprehensive study of fluid spheres made of anisotropic matter was presented in [62] in Newtonian as well as relativistic gravity. In the context of Einstein’s General Relativity, anisotropic fluids have been extensively studied, with particular emphasis on stellar modeling and their impact on structural properties of compact objects [63, 64].
@@
This work demonstrates the necessary steps to derive a consistent solution in the context of $f(Q)$ gravity, using a choice of a connection that is compatible with any $f(Q)$ theory and allows for the satisfaction of the connection equation without the generation of additional constraints on the metric function outside of the metric equations. The primary objective of the article is to establish the mathematical and conceptual consistency of stellar solutions in $f(Q)$ gravity, rather than to propose a definitive astrophysical model. For this reason, in the example we present, we opt for the derivation of an exact solution by imposing a function $f(Q)$ of a power-law form and adopting a specific ansatz for the metric.
@@ III.1 Relativistic stars within General Relativity
When seeking interior solutions ($0\leq r\leq\mathcal{R}$, $\mathcal{R}$ being the radius of the star), the most general form of a static, spherically symmetric geometry in Schwarzschild-like coordinates, $\{t,r,\theta,\phi\}$, is given by
@@
$p_{r}(\mathcal{R})=0,\;\;\;\;m(\mathcal{R})=M,\;\;\;\;e^{A(\mathcal{R})}=1-2\frac{M}{\mathcal{R}},$ (25)
with $M$ being the stellar mass, taking into account that the exterior vacuum solution ($r>\mathcal{R}$) is given by the Schwarzschild geometry [60]
@@
$A(r)=\ln\left(1-\frac{2M}{\mathcal{R}}\right)+2\int^{r}_{\mathcal{R}}dx\>\frac{m(x)+4\pi x^{3}p_{r}(x)}{x^{2}\left(1-2m(x)/x\right)}.$ (27)
@@ III.2 Exact analytic solution within $f(Q)$ gravity
The theory is characterized by two basic fields, the metric and the connection. For an excellent review of the geometric aspects of theories with nonmetricity we refer the interested reader to [28] and references therein. For the former we adopt the static and spherically symmetric line element as seen in Eq. (17). In what regards the connection, it is well known [77] that for a flat and symmetric connection there can always be found a coordinate system where its components are zero, i.e. $\Gamma_{~\mu\nu}^{\alpha}=0$. This is referred to in the literature as the coincident gauge. However, we need to be careful when we impose such a condition. When assuming a particular class of metrics, like the one we read from (17), we need to consider that we have already fixed the gauge by choosing a particular coordinate system. The latter may happen to be incompatible with having $\Gamma_{~\mu\nu}^{\alpha}=0$.
In these cases, a guide in order to write connections which are compatible with the field equations, is given by requiring that the connection shares the symmetries of the metric, i.e. the Killing vectors
$\begin{split}&\xi_{1}=\frac{\partial}{\partial\phi}~,~\xi_{2}=\cos\phi\frac{\partial}{\partial\theta}-\sin\phi\cot{\theta}\frac{\partial}{\partial\phi}\\ &\xi_{3}=\sin\phi\frac{\partial}{\partial\theta}+\cos\phi\cot{\theta}\frac{\partial}{\partial\phi}~,~\xi_{4}=\partial_{t}.\end{split}$ (28)
The first three are the generators of spherical symmetry and the fourth imposes staticity. We demand that the above isometries, satisfying $\mathcal{L}_{\xi_{i}}g_{\mu\nu}=0$, where $\mathcal{L}$ is the Lie-derivative and i=1,,4i=1,...,4, leave invariant the connection as well
$\mathcal{L}_{\xi_{i}}\Gamma_{~\mu\nu}^{\alpha}=0,\quad i=1,...,4.$ (29)
The resulting system of partial differential equations has two general families of solutions [78, 47]. They both play an an important role in the dynamics as they introduce additional degrees of freedom [37]. The first family of solutions, leads to off-diagonal components in the field equations (10) for the metric, which can only be eliminated by setting Q=Q=const. or $f(Q)\sim Q$. Alternatively, one can keep the off-diagonal terms and consider a non-diagonal energy momentum tensor for the matter.
For the second family of connections, the relative non-diagonal terms can be removed without requiring a constraint on the theory $f(Q)$ or on $Q$ itself. It is this second connection which we choose in this work. Its non-zero components are [78, 47, 50]:
@@
Nonmetricity theories are based on two basic fields, the metric $g_{\mu\nu}$ and the connection $\Gamma_{\;\mu\nu}^{\kappa}$. The set of components (30) forms the initial ansatz for the connection $\Gamma_{\;\mu\nu}^{\kappa}$ in the same manner line element (17) constitutes the ansatz for the metric. For this choice of a connection, and for metric (17), the off-diagonal terms of the field equations (10) are eliminated if we further choose $c_{1}=c_{2}=0$.
@@
satisfies equation (32) irrespectively of the $f(Q)$ function and of the $A(r)$ and $B(r)$ appearing in the metric. This connection has been also utilized in the context of scalar nonmetricity to derive black hole solutions [22]. We should underline here that the $\gamma_{2}$ of (34) is not the general solution of the connection equation (32). However, it has the advantage of being the only choice satisfying Eq. (32) irrespectively of the $f(Q)$ function under consideration. Any other solution of (32), including the most general configuration of (30) with non-zero $c_{1}$ and $c_{2}$, would be applicable to a particular $f(Q)$ function. The utilization of such a solution allows for the direct comparison among different $f(Q)$ theories by maintaining a single ansatz for both fundamental fields, i.e. the metric tensor and the connection.
A brief comment is in order at this point. Many works in the literature, attempting to study stellar configurations in $f(Q)$ gravity, either incorporate an over-restricting procedure or include the serious error of disregarding the equation of the connection. In some cases, the coincident gauge is adopted, which however is incompatible with line element (17) in spherical coordinates and thus overrestricts the equations; leading only to solutions of the linear $f(Q)\sim Q$ theory, which is however equivalent to GR [79, 80, 81, 82, 83]. In other cases, even though a connection compatible with spherical symmetry is considered, it is usually not checked whether the equation of motion for the connection (12) is satisfied. To make this problem more explicit; a common choice of a connection used in the literature [84, 85, 86, 87, 88] is the following
@@
Remember that $Q$ here is given by (36), and for a given $f(Q)$ theory this is going to be purely a relation between the functions of the metric, $A$ and $B$. The above equation is trivialized if Q=Q=const. or $f(Q)\sim Q$, which, as is known, describe the same solutions as in the case of GR with or without cosmological constant. Thus, in the generic setting of a non-trivial $f(Q)$ theory, it is not sufficient for the description of the star solution to just solve algebraically (31) with respect to the energy density and the pressures. The constraint set by the field equation of the connection needs to be satisfied. Note that it is not enough to satisfy the analogue of the Tolman-Oppenheimer-Volkoff (TOV) equation. The $\nabla_{\mu}T^{\mu\nu}=0$ condition on the matter content together with the field equations for the metric (31) are not sufficient conditions. This is because, as we previously mentioned, the covariant derivative of the gravitational tensor $E_{\mu\nu}$ of Eq. (13), unlike what happens in GR, is not identically zero. The $\nabla_{\mu}E^{\mu\nu}=0$ equation holds only if the equation of the connection (32) is satisfied [74]. In order for the latter to happen for a theory dynamically different from GR, we need to find an appropriate $\gamma_{2}(r)$ function that solves the connection equation. The importance of using a connection compatible with the field equations has also been stressed recently in [57].
In our case, the $\gamma_{2}(r)$ we derived in (34) serves to satisfy the field equation for the connection (32), and thus the constraint in the metric functions is overall avoided. What is more, (34) satisfies the relevant equation for any $f(Q)$ theory. In this respect, the $\gamma_{2,\text{int}}=re^{-\frac{B}{2}}$ forms an excellent choice for the connection in what regards the interior solution for the star. For the exterior vacuum solution we adopt a different path, which we describe in detail in the next section. It consists of adopting an exterior solution, $\gamma_{2,\text{ext}}$, for which the nonmetricity scalar becomes constant. Thus, we achieve for the inside, a consistent description of modified gravitational dynamics due to a dynamical nonmetricity, while in the exterior a known vacuum solution compatible with GR plus a cosmological constant. In the same manner that one connects two distinct solutions for the interior (with matter) and for the exterior (vacuum) for the metric, we additionally connect here the two distinct solutions for the other basic field of the theory, the connection.
@@ III.3 Exterior solution
where the effective cosmological constant $\Lambda_{\text{eff}}$ is given by (15). Note that the resulting Schwarzschild-De Sitter geometry forms the unique type of $f(Q)$ vacuum black-hole solution with reciprocal metric functions $g_{tt}=-1/g_{rr}$ [54].
@@
Schwarzschild-(Anti-)De Sitter spacetimes in alternative theories of gravity often arise as exterior black hole solutions with constant curvature; for example in $f(R)$ theories of gravity. In our case it is the constant $Q$ and not the constant $R$ that forms the fundamental scalar of the theory. Note that despite of having a constant nonmetricity outside the star, the interior dynamics is still governed by the non-linear $f(Q)$ terms, which dictate the stellar structure and stability limits in a way distinct from that of Einstein’s theory.
@@ IV The model under consideration
Our following analysis is split into two parts: First, we calculate in a concise manner the exact analytic solution and all the quantities of interest, such as: pressure, energy density, speed of sound, etc. Then, we proceed to investigate what kind of stars this solution represents and how they compare to the ones in the context of GR.
IV.1 The exact interior solution
@@
where $\alpha$, $Q_{0}$ and $\nu$ are constants. Even though our main objective is to use this theory to model a basic example as part of a concise mathematical procedure, this choice is not without physical justification. We already know that the linear theory is dynamically equivalent to GR, and we are also aware that deviations from Einstein’s theory cannot be arbitrarily large if we are to satisfy well-established solar system tests. The addition of a polynomial term with a parameter constrained to be small is the simplest modification to this theory. At the same time, by adopting the particular value $\nu=2$ for the exponent (we present for this value the basic graphs in the analysis that follows), the resulting expression consists of the first two terms of a Taylor expansion of an arbitrary $f(Q)$ theory in the region of a small nonmetricity, $Q<<1$. Thus, this choice of $f(Q)$ function can serve to draw conclusions valid as the asymptotic limit of more general theories.
The ansatz we follow for the metric functions, $A(r)$ and $B(r)$ is the following [89]
@@
with $C\in\mathbb{R}$ and $r_{1}$, $r_{2}$ positive constants. The corresponding geometry describes a singular-free physical solution. Since the solution is based on a metric ansatz without assuming a certain equation-of-state, it should be understood as an effective model rather than a microphysically realistic description.
@@
Using Eq. (46) in the expressions (43)-(45), and considering theories $f(Q)$ for which $\nu>1$ in (40), it is straightforward to demonstrate that the relevant physical quantities assume a finite value at $r=0$. The nonmetricity scalar itself becomes zero as $r\rightarrow 0$. In the Taylor expansion, there always exists a zero-th order term, while the rest that follow are in positive powers of $r$. Specifically, for $\nu\geq 2$ the leading $r$-dependent terms are quadratic:
@@
$c_{s,r}^{2}=\frac{p_{r}^{\prime}}{\rho^{\prime}},\quad c_{s,t}^{2}=\frac{p_{t}^{\prime}}{\rho^{\prime}}$ (50)
respectively, we obtain at the limit $r\rightarrow 0$
$\lim\limits_{r\to 0}c_{s,r}^{2}=\frac{\left(4r_{2}^{2}-r_{1}^{2}\right)\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+4\alpha r_{2}^{2}\right)}{5r_{1}^{2}\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+8\alpha r_{2}^{2}\right)},$ (51)
$\lim\limits_{r\to 0}c_{s,t}^{2}=\frac{4\alpha\left(r_{1}^{4}-6r_{1}^{2}r_{2}^{2}+8r_{2}^{4}\right)-2Q_{0}r_{2}^{2}\left(r_{1}^{4}-3r_{1}^{2}r_{2}^{2}+r_{2}^{4}\right)}{5r_{1}^{2}\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+8\alpha r_{2}^{2}\right)}$ (52)
@@
$\lim\limits_{r\to 0}c_{s,r}^{2}=\frac{4r_{2}^{2}}{5r_{1}^{2}}-\frac{1}{5},$ (53)
$\lim\limits_{r\to 0}c_{s,t}^{2}=\frac{2\left(3r_{1}^{2}r_{2}^{2}-r_{1}^{4}-r_{2}^{4}\right)}{5r_{1}^{4}}.$ (54)
IV.2 Physical relevance of the solution
In what follows, we adopt specific values for the relative parameters for the model, in order to investigate the behavior of the solution and what type of stars it predicts. In each case, we make the relative comparison to GR demonstrating under which different conditions the same metric potentials are produced. We investigate the impact of the coupling $\alpha$, which measures the deviation of GR, on the structural properties of the stars. We also check whether or not well established criteria are met.